OpenStudy (anonymous):

factoring:

OpenStudy (anonymous):

describe and correct the error made in factoring completely. 4x^4+12x^3+8x^2+24x=4(x^4+3x^3+2x^2+6x) =4[x^3(x+3)+2x(x+3)] =4(x^3+2x)(x+3)

OpenStudy (cwrw238):

the only thing I can see is the omission of the square brackets in the last step

OpenStudy (anonymous):

its number 35

OpenStudy (cwrw238):

i can see that document i'm sfraid

OpenStudy (anonymous):

cant see it?

OpenStudy (anonymous):

well how am i supposed to show it to you then?

OpenStudy (cwrw238):

ok i've managed to expand it i cant see anything else wrong - but thats not really an error its the same without square brackets

OpenStudy (anonymous):

are you sure?

OpenStudy (cwrw238):

yes - perhaps its a catch question

OpenStudy (anonymous):

oh okay well can i ask you other questions about factoring?

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

why can you factor the same polynomial using different pairs of terms?

OpenStudy (anonymous):

describe how to factor the expression 6x^5+4x^4+12x^3+8x^2+9x+6.

OpenStudy (cwrw238):

try to do this by grouping 9x + 6 = 3 (3x + 2) 6x^5 + 4x^4 = 2x^4(3x + 2) and 12x^3 + 8x^2 = 4x^2 (3x + 2) so we have 3x + 2 common to all 3 so we can write (3x + 2)( 2x^4 + 4x^2 + 3)

OpenStudy (cwrw238):

now we have a trinomial in the second bracket which might be factorable..

OpenStudy (cwrw238):

no its not so thats the final answer

OpenStudy (anonymous):

ok?

OpenStudy (cwrw238):

yes I was looking at a binomial of the form (2x^2 + a)(x^2 + b) for the second bracket but the numbers will not fit

OpenStudy (anonymous):

your confusing me?

OpenStudy (cwrw238):

compare it with a trinomial which factorises 2x^2 +5x + 3) = 2x^2 + 2x + 3x + 3 = 2x(x + 1) + 3(x + 1) = (2x + 3)((x + 1)

OpenStudy (cwrw238):

theers no way can you get whole numbers to get 4x as the middle term